Show how both sides equal.
sin 45 degrees= Square root 1-cos 90 degrees/2.
PLEASE HELP

Answer by eperette(173)

things you should know:
sin45= sqr(2)/2
cos90= 0

sin45 = sqr(1-cos90/2)
sqr(2)/2 = sqr (1-0/2)
sqr(2)/2 = sqr (1/2)
sqr(2)/2 = sqr(1)/sqr(2)
sqr(2)/2 = 1/sqr(2)
sqr(2)/2 = 1/sqr(2) * sqr(2)/sqr(2)
sqr(2)/2 = sqr(2)/2